OpenStudy (anonymous):

5. Show all steps used to solve the problem. One point for steps and one point for solution. Be sure that you check your solution for correctness. If 3a - 9 = 6, what is the value of 4a - 1?

OpenStudy (anonymous):

igonore the 1 point for each coreect solution

OpenStudy (anonymous):

3a=6+9 3a=15 a=15/3 a=5 Substitute a in 4a-1 4(5)-1 20-1 =19

OpenStudy (anonymous):

i dont get that...

OpenStudy (anonymous):

the 3a part is final i doesnt need to be solved

OpenStudy (anonymous):

i didn't get u sry what u didn't get